OpenStudy (anonymous):

HELP I GOT THE WRONG ANSWER 5. What is the solution to the following equation 4 - 2y = -8 A. y = 2 B. y = -6 C. y = -2 D. y = 6 I choose answer C but my teacher said it was wrong.

OpenStudy (carniel):

-8 -4 = -12/-2 = 6 D.

OpenStudy (anonymous):

@Carniel we have different answers i guess

OpenStudy (carniel):

You subtract 4 from both sides which gives you -12 then divide by -2

OpenStudy (anonymous):

@Carniel wait, nvm, you're right lol I made a mistake

OpenStudy (carniel):

Negative divided by a negative is a positive

OpenStudy (anonymous):

4 - 2y = -8 -2y=-8-4 -2y=-12 -12/-2y y=6 Which is D :)